# Data Fusion for Heterogeneous Treatment Effect Estimation with Multi-Task Gaussian Processes

### v1

@@ 4 Experiments
We investigated Causal-ICM’s performance through multiple simulation studies and a comprehensive analysis of Real-World Data (RWD) sourced from the Tennessee STAR study (Achilles et al., 2008). The code to reproduce the experiments is available on Github  [^1] https://github.com/EvanDimitriou/CausalICM. The optimal value of $\rho$ is chosen data-adaptively in all cases (Section 3.4.2). The evaluation of various models and their comparison was conducted using the RMSE where we average over the observational study’s covariate distribution, which we treat as the target population. We compared Causal-ICM to the T-learner approach with GPs as regressors trained on each study separately and the 2-step method of Kallus et al. (2018), which debiases observational data using experimental data through a low-complexity bias function, with either Random Forest (RF) or GPs as base learners. Additionally, we assessed the coverage of the credible intervals for the CATE produced by Causal-ICM across different simulation settings and compared with the T-learner approaches. In the multivariate covariate case, we consider the pointwise coverage of the CATE averaged over the observational study’s covariate distribution. We used GPy (2012) to train the ICMs, with $\rho$ tuned as described in Section 3.4.2 and remaining hyperparameters estimated via marginal likelihood maximisation. All experiments were run on a single Macbook Pro (M2) with 16Gb RAM.
@@ 4.1 Simulation Studies
We showcase the performance of Causal-ICM in the main text using two distinct simulation studies (univariate and multivariate). Further simulation studies are detailed in Appendix A.8, illustrating the competitive performance of Causal-ICM in a wide range of settings. In all simulations, the sample size is roughly between 200-300 for the RCT (due to random selection) and 1000 for the observational study.
@@
In the univariate simulation setting, we assume a continuous baseline covariate $X\sim U\left[-2,2\right]$. The values of $X$ control the trial participation probability, where $S\sim\text{Bernoulli}\left(p_{S}\right)$ for $p_{S}=\frac{\text{exp}\left(-3-3X\right)}{1+\text{exp}\left(-3-3X\right)}$. Here, patients with low $X$ values have higher probability of being assigned in the trial. After selecting all trial participants, we assign them randomly to treatment with $A\sim\text{Bernoulli}\left(0.5\right)$. The potential outcomes are generated as $Y\left(a\right)=a\tau\left(X\right)+X^{2}-1+\epsilon$ for $a=0,1$, where $\tau\left(X\right)=1+X+X^{2}$ is the CATE and $\epsilon\sim\mathcal{N}\left(0,1\right)$. For the observational study we have $X\sim U\left[-2,2\right]$ and the treatment is generated according to the model $A\sim\text{Bernoulli}\left(e\left(X\right)\right)$ where $\text{logit}\left(e\left(X\right)\right)=-X$. Similarly to the trial, the potential outcomes are $Y\left(a\right)=a\tau\left(X\right)+X^{2}-1+U+\epsilon$ for $a=0,1$. Here, $U$ is the hidden confounder which is generated from $U\sim\mathcal{N}\left(\left(2A-1\right)\text{sin}\left(X-1\right),1\right)$, yielding a nonlinear confounding effect.
[…]
The multivariate simulation study was generated similarly with 5 covariates, with some confounding the study participation or the treatment allocation mechanism. More details can be found in Appendix (A.8.2). In the multivariate case (where $\rho=0.3$ was selected), Causal-ICM also produced the lowest average RMSE value ($2.46\pm 0.15$), followed by the T-learner trained on the experimental data ($3.08\pm 0.20$), 2-step method of Kallus et al. (2018) ($3.16\pm 0.16$) and last by the observational-only T-learner ($3.46\pm 0.05$). Here, we report the mean $\pm$ 1 standard error, and results for the bias can be found in Appendix A.8.2. Causal-ICM was again conservative with higher than nominal level coverage (Table 2). The lowest average coverage is again attained by the observational-only T-learner. We note that TGP-exp achieves slightly narrower interval lengths. This is likely due to kernel hyperparameter optimization variability: we expect the Causal-ICM intervals to be narrower when the kernel hyperparameter values are chosen to be the same as TGP-exp.
@@ 4.2 Real-World Data Analysis
The Tennessee Student/Teacher Achievement Ratio (STAR) (Achilles et al., 2008) randomised experiment initiated in 1985 studied the impact of class size on student outcomes through standardized test scores from first to third grade. We focus on two ‘treatments’, here corresponding to two experimental conditions: small class size and regular class size. We followed a similar approach to Kallus et al. (2018) and used the real data to produce a ‘confounded’ dataset, corresponding to the observational sample, and a smaller unconfounded dataset, which comprised our RCT data. After removing subjects with missing treatment or outcome values, we were left with a randomised sample of 4218 students. Baseline covariates included gender, race, birth month, birth year, free lunch eligibility and teacher ID. To obtain the ‘experimental’ dataset we sampled randomly only from rural or inner city students. To form the confounded observational dataset, we aggregated all control cases not included in the experimental data. Additionally, for those undergoing treatment, we drew a sample with a down-weighting mechanism applied to individuals whose outcomes fell below the $30^{th}$ percentile. After preprocessing, the unconfounded dataset included 422 students, while the confounded dataset included 2593. Further, 379 students were kept as a validation set.
@@
In order to compare Causal-ICM with existing methods, we require a notion of ground truth. Given the randomised nature of the study, we assume that an unbiased estimate of the CATE can be obtained using a doubly robust estimator, which we then treat as the ground truth. To obtain this, we estimated the propensity score and conditional expectation models using the full dataset, prior to splitting it into the confounded and unconfounded samples (Saito and Yasui, 2020). We note that, in the absence of the true CATE, other performance metrics adapted to the counterfactual nature of the evaluation are also available (Alaa and Van Der Schaar, 2019; Boyer et al., 2023). In the RWD analysis, Causal-ICM achieved the best performance in terms of RMSE followed by the 2-step approach of Kallus et al. (2018) with RF as base learners (Table 2).
Table 1: Multivariate simulation setting:
Average coverage and interval lengths
Coverage Length
Nominal $90\%$ $95\%$ $90\%$ $95\%$
Causal-ICM ($\rho=0.3$) $98.5\%$ $100\%$ 4.72 5.62
TGP-obs $69\%$ $78\%$ 1.01 1.20
TGP-exp $93\%$ $96.5\%$ 4.35 5.18
Table 2: RWD Analysis:
Model comparison
RMSE
TGP-obs 57.89
TGP-exp 43.17
2-step GP 41.62
2-step RF 39.07
Causal-ICM ($\rho=0.1$) 37.86
[…]
@@ 5 Discussion
We have developed Causal-ICM, a rank-2 ICM to combine observational and experimental data to estimate treatment effects for a target population, where the hyperparameter $\rho$ controls the degree of borrowing in an interpretable way. By providing accurate estimates of heterogeneous treatment effects, Causal-ICM has the potential to inform more personalised decision making in medicine. Moreover, by highlighting parts of the population for which treatment effects are highly uncertain, our methodology may guide the design of future trials to improve their representativeness and reduce this uncertainty to a satisfactory level. Although extrapolation is generally challenging, the ICM appears to extrapolate successfully with calibrated uncertainty quantification, achieving reasonable levels of coverage. Nonetheless, it is challenging to guarantee frequentist coverage for Bayesian nonparametric methods, which is one of the limitations of our method. Another limitation is the sensitivity of uncertainty estimates to the kernel hyperparameter optimization, which is an intrinsic challenge of GP s. Future directions or research include the consideration of multiple data sources or multiple treatment arms, imposing a prior distribution on the parameter $\rho$, and incorporating the propensity score into our model. Investigating the partial identification of causal effects under Causal-ICM, especially in regions of extrapolation, is also an important area for future work.
Appendix A Appendix
A.1 Identifiability assumptions
We outline the identifiability assumptions for the CATE in the RCT population below.
  1. 1.
    Consistency of counterfactual outcomes: If $A=a$ then $Y^{a}=Y$.
    The intervention is well defined, there are no hidden versions of the treatment. This assumption can be supported by the study design, i.e. careful implementation of standardized treatment protocols.
  2. 2.
    No interference between individuals of the treatment groups, i.e. one subjects treatment cannot affect another subject’s outcome.
  3. 3.
    No hidden confounding in the trial: $Y^{a}\perp\!\!\!\perp A|\left(\mathbf{x},S=1\right)$
    This assumption is supported by randomization.
  4. 4.
    Positivity of treatment in the trial: for every covariate level $\mathbf{x}$ with positive density in the trial (i.e. $\mathbf{x}\in\mathcal{X}^{\text{e}}$), the probability of being assigned each of the randomized treatments is positive, i.e. $\mathbb{P}\left[A=a|\textbf{x},S=1\right]>0$
A.2 Illustrative example
@@
Figure 2: Comparative illustration between Independent GPs and Multitask GP. Here, $A$ represents distinct treatment groups, $Y$ corresponds to the outcome, and $X$ denotes baseline covariates.
A.3 Derivation of Posterior Distribution
@@
$\mathbb{E}\left[y^{\text{e}}_{i}y^{\text{e}}_{j}\right]$ =$=\begin{cases}\beta^{\text{e}}k(\mathbf{x}^{\text{e}}_{i},\mathbf{x}^{\text{e}}_{j})+\sigma^{2}\quad&\text{if }i=j\\ \beta^{\text{e}}k(\mathbf{x}^{\text{e}}_{i},\mathbf{x}^{\text{e}}_{j})\quad&\text{otherwise}\end{cases}$
@@
$\mathbb{E}\left[y^{\text{o}}_{i}y^{\text{o}}_{j}\right]$ =$=\begin{cases}\beta^{\text{o}}k(\mathbf{x}^{\text{o}}_{i},\mathbf{x}^{\text{o}}_{j})+\sigma^{2}\quad&\text{if }i=j\\ \beta^{\text{o}}k(\mathbf{x}^{\text{o}}_{i},\mathbf{x}^{\text{o}}_{j})\quad&\text{otherwise}\end{cases}$
@@
$\begin{bmatrix}\mathbf{y}\\ \mathbf{f}(\mathbf{x}_{*})\end{bmatrix}\sim\mathcal{N}\left(\mathbf{0},\begin{bmatrix}K(X,X)+\sigma^{2}I&K(X,\mathbf{x}_{*})\\ K(\mathbf{x}_{*},X)&K(\mathbf{x}_{*},\mathbf{x}_{*})\end{bmatrix}\right)$
@@
A.4 Proof of Proposition 3.1
@@
$B^{\prime}=\begin{bmatrix}(\beta^{\text{e}})^{2}&\beta^{\text{e}}\beta^{\text{eo}}\\ \beta^{\text{e}}\beta^{\text{eo}}&(\beta^{\text{eo}})^{2}\end{bmatrix},\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ k^{\prime}(\mathbf{x}_{*},\mathbf{x}_{*})=k(\mathbf{x}_{*},X^{\text{e}})^{T}{\Sigma_{\text{e}}}^{-1}k(\mathbf{x}_{*},X^{\text{e}})$
@@
A.5 Interpreting $\rho$
Proposition A.1.
We have $\rho=1$ if and only if
Proof.
It is easier to work with $1-\rho^{2}$, where plugging in the values from (1) gives
The denominator is positive and finite, so the numerator is thus 00 if and only if $\alpha_{1}^{\text{e}}\alpha_{2}^{\text{o}}=\alpha_{1}^{\text{o}}\alpha_{2}^{\text{e}}$. ∎
A.6 Confounding function

### v2

@@ 4 Experiments
We investigated Causal-ICM’s performance through multiple simulation studies and a comprehensive analysis of Real-World Data (RWD) sourced from the Tennessee STAR study (Achilles et al., 2008). The code will be made available online. The optimal value of $\rho$ is chosen data-adaptively in all cases (Section 3.4.2). Model evaluation and comparison were based on the RMSE, averaged over the covariate distribution of the observational study.
We compared Causal-ICM against several benchmarks: (i) a T-learner with GPs regressors trained separately on each study, (ii) the two-step method of Kallus et al. (2018), which debiases observational data via a low-complexity bias function with GPs as base learners, (iii) the integrative estimator of Yang et al. (2022b), which assumes a non-linear CATE and linear confounding effect, (iv) the power likelihood method of Lin and Evans (2023), and (v) the test-then-pool approach of Yang et al. (2022a). For Yang et al. (2022b), we used the default settings recommended in their paper, while for Lin and Evans (2023), we selected the power parameter $\eta$ by maximizing the Expected Log Pointwise Predictive Density (ELPD), which in both cases resulted in $\eta=0$ (no pooling). Finally, we assessed the coverage of Causal-ICM’s CATE credible intervals across simulation scenarios, comparing them with those from the T-learner approaches.
We used GPy (2012) to train the ICMs, with $\rho$ tuned as described in Section 3.4.2 and remaining hyperparameters estimated via marginal likelihood maximisation. All experiments were run on a single Macbook Pro (M2) with 16GB RAM.
@@ 4.1 Simulation Studies
We showcase the performance of Causal-ICM in the main text using two distinct simulation studies. Further simulation studies, both univariate and multivariate, are detailed in Appendix A.8, illustrating the competitive performance of Causal-ICM in a wide range of settings, including both linear and non-linear scenarios for the potential outcomes, CATE and confounding function. In all simulations, the sample size is roughly between 200-300 for the RCT and 1000 for the observational study.
@@
In the first simulation setting, we assume a continuous baseline covariate $X\sim\text{Unif}\left[-2,2\right]$. The values of $X$ control the trial participation probability, where $S\sim\text{Bernoulli}\left(p_{S}\right)$ for $p_{S}={\text{exp}\left(-3-3X\right)}/{(1+\text{exp}\left(-3-3X\right)})$. Here, patients with smaller $X$ values have higher probability of being assigned in the trial. After selecting all trial participants, we assign them randomly to treatment with $A\sim\text{Bernoulli}\left(0.5\right)$. The observed outcomes are generated as $Y=A\tau\left(X\right)+X+\epsilon$ for $A\in\{0,1\}$, where $\tau\left(X\right)=1+X$ is the CATE and $\epsilon\sim\mathcal{N}\left(0,1\right)$. For the observational study we have $X\sim\text{Unif}\left[-2,2\right]$ and the treatment is generated according to $A\sim\text{Bernoulli}\left(e\left(X\right)\right)$ where $\text{logit}\left(e\left(X\right)\right)=-X$. Similarly to the trial, the observed outcomes are $Y=A\tau\left(X\right)+X+U+\epsilon$ for $A\in\{0,1\}$. Here, $U$ is the hidden confounder which is generated from $U\sim\mathcal{N}\left(\left(2A-1\right)X,1\right)$, yielding a linear confounding effect $\eta(X)=2X$. In the second simulation, we change the potential outcomes, the CATE and the confounding effect to be non-linear as follows: $Y=A\tau(X)+X^{2}-1+U+\epsilon$, where $\tau(X)=1+X+X^{2}$ and $U\sim\mathcal{N}((2A-1)\sin(X-1),1)$, yielding $\eta(X)=2\text{sin}(X-1)$. All other variables remain the same.
Under a linear CATE and a linear unobserved confounding effect, Causal-ICM (with $\rho=0.1$) exhibited comparable performance to the integrative HTE estimator (Yang et al. (2022b)) and the Elastic HTE estimator of Yang et al. (2022a), and superior performance to all other techniques (Fig 1 left). In the non-linear setting (both CATE and unobserved confounding effect), Causal-ICM outperformed all other methods yielding the lowest RMSE (Fig 1 right). These results reflect the limitations of the experimental grounding, the Integrative HTE and the Elastic HTE methods. Each of these are designed for linear treatment or confounding effect, resulting in inflated RMSE values in scenarios where these assumptions do not hold. In the multivariate setting Causal-ICM performed equally good or better than most methods, being outperformed by parametric approaches only when the model was correctly specified (see Appendix G.2 Fig 4).
Regarding uncertainty quantification, our approach generated 95% credible intervals with reasonable, though slightly conservative, coverage close to the nominal level in both simulation settings (Fig 2). The observational-only T-learner had the lowest coverage across the baseline covariate support, which is expected since its estimates are biased but exhibit lower uncertainty due to the larger dataset. This is exactly the behaviour that Causal-ICM is robust to, as formalized in Proposition 3.1: the large sample size of the observational study does not overwhelm the Causal-ICM’s estimates. See Figure 5 in Appendix H for an illustrative example of extrapolation and uncertainty quantification with Causal-ICM in the second simulation setting. Additionally, the Causal-ICM is robust against the kernel choice (see Appendix I Table 3), the observational sample size (see Appendix I Table 4), and the degree of overlap between the two studies (see Appendix I Table 5), yielding similar RMSE values under all these different scenarios. Finally, the choice of $\rho$ plays an important role in controlling the influence of the confounded observational data (see Appendix I Table 2).
Figure 1: Simulation results over 100 simulated datasets. Left: Simulation setting 1: linear CATE and unobserved confounding, Right: Non-linear CATE and unobserved confounding
Figure 2: 95% conditional coverage over the covariate distribution of the observational study for the two simulation studies. The black solid line indicates the nominal level (95%). Left: First simulation setting (linear CATE and confounding), Right: Second simulation setting (non-linear CATE and confounding). The GP (obs) method exhibited 0% coverage in most of the support covariate distribution.
@@ 4.2 Real-World Data Analysis
The Tennessee Student/Teacher Achievement Ratio (STAR) (Achilles et al., 2008) randomised experiment initiated in 1985 studied the impact of class size on student outcomes through standardized test scores from first to third grade. We focus on two ‘treatments’, here corresponding to two experimental conditions: small class size and regular class size. We followed a similar approach to Kallus et al. (2018) and used the real data to produce a ‘confounded’ dataset, corresponding to the observational sample, and a smaller unconfounded dataset, which comprised our RCT data. After removing subjects with missing treatment or outcome values, we were left with a randomised sample of 4218 students. Baseline covariates included gender, race, birth month, birth year, free lunch eligibility and teacher ID. To obtain the ‘experimental’ dataset we sampled randomly only from rural or inner city students. To form the confounded observational dataset, we aggregated all control cases not included in the experimental data. Additionally, for those undergoing treatment, we drew a sample with a down-weighting mechanism applied to individuals whose outcomes fell below the $30^{th}$ percentile. After preprocessing, the unconfounded dataset included 422 students, while the confounded dataset included 2593. Further, 379 students were kept as a validation set. We compared Causal-ICM (optimal $\rho=0.4$) with a GP based T-learner trained either on the experimental or the observational data,the experimental grounding method of Kallus et al. (2018) either with Random Forest or GP as base learners and the Integrative HTE method of Yang et al. (2022b). We were unable to include the Elastic HTE estimator (Yang et al. (2023)) or the power likelihood approach (Lin et al. (2025a)), as both methods encountered convergence issues due to the datasets’ high dimensionality and the limited variability in their categorical covariates.
@@
In order to compare Causal-ICM with existing methods, we require a notion of ground truth. Given the randomised nature of the study, we assume that an unbiased estimate of the CATE can be obtained using a doubly robust estimator, which we then treat as the ground truth. To obtain this, we estimated the propensity score and conditional expectation models using the full dataset, prior to splitting it into the confounded and unconfounded samples (Saito and Yasui, 2020). We note that, in the absence of the true CATE, other performance metrics adapted to the counterfactual nature of the evaluation are also available (Alaa and Van Der Schaar, 2019; Boyer et al., 2023).
The results are summarised in Table 1 below. Causal-ICM again performed highly competitively in terms of RMSE, on a par with the experimental grounding method with Random Forests as base learners.
Table 1: RMSE values for the RWD example. Causal-ICM: our method; GP (exp): T-learner with GPs as base learners trained in the unconfounded; GP (obs): T-learner with GPs as base learners trained in the confounded data; Experimental grounding (GP): the method proposed by Kallus et al. (2018) with GPs as base learners; Experimental grounding (RF): the method proposed by Kallus et al. (2018) with Random Forest as base learners; Integrative HTE: the method proposed by Yang et al. (2022b)
Method RMSE     Method RMSE
  Causal-ICM ($\rho=0.4$) 6.29     Experimental grounding (GP) 6.38
GP (exp) 6.36     Experimental grounding (RF) 6.27
GP (obs) 7.47     Integrative HTE 11.84
@@ 5 Discussion
We have developed Causal-ICM, a rank-2 ICM to combine observational and experimental data to estimate treatment effects for a target population, where the hyperparameter $\rho$ controls the degree of borrowing in an interpretable way. In contrast to existing methods, Causal-ICM is highly flexible and does not make any assumptions about the functional form of the CATE or the effect of unobserved confounding. By providing accurate estimates of heterogeneous treatment effects, Causal-ICM has the potential to inform more personalised medical decision making. Moreover, unlike other data fusion methods, Causal-ICM is able to provide uncertainty quantification. By identifying regions of the population where treatment effects remain highly uncertain, our approach can help guide the design of future trials to improve representativeness and reduce uncertainty to satisfactory levels (Yang et al. (2025)).
Despite the inherent difficulty of extrapolation, Causal-ICM appears to extrapolate well, achieving reliable levels of coverage. In general, it is challenging to ensure frequentist coverage for Bayesian nonparametric methods, which is a limitation of our approach. Another limitation is the sensitivity of uncertainty estimates to the kernel choice and the hyperparameter optimization, which is an intrinsic challenge of GPs. Future research directions for Causal-ICM include extensions to multiple data sources or treatment arms, placing a prior on the parameter $\rho$, and incorporating the propensity score into the model. Investigating partial identification of causal effects, especially in regions of extrapolation, is also important Pfister and Bühlmann (2024). Methodological extensions such as allowing distinct kernels for the latent functions, accommodating non-Gaussian likelihoods, and improving computational scalability via sparse GPs or inducing points represent additional promising directions (see Appendix J).
Appendix A Identification of CATE in the RCT
Below we show that the CATE can be identified using observed data in the RCT.
$\mathbb{E}[Y^{1}-Y^{0}\mid\mathbf{X}_{\tau}]$ $=\mathbb{E}\!\left[\mathbb{E}[Y^{1}-Y^{0}\mid\mathbf{X}_{\tau}\cup\mathbf{X}_{S}]\mid\mathbf{X}_{\tau}\right]$
(tower property)
$=\mathbb{E}\!\left[\mathbb{E}[Y^{1}-Y^{0}\mid\mathbf{X}_{\tau}\cup\mathbf{X}_{S},S=\mathrm{e}]\mid\mathbf{X}_{\tau}\right]$
(Assumptions A4)
$=\mathbb{E}\!\left[\mathbb{E}[Y^{1}\mid A=1,\mathbf{X}_{\tau}\cup\mathbf{X}_{S},S=\mathrm{e}]-\mathbb{E}[Y^{0}\mid A=0,\mathbf{X}_{\tau}\cup\mathbf{X}_{S},S=\mathrm{e}]\mid\mathbf{X}_{\tau}\right]$
(Assumptions A2)
$=\mathbb{E}\!\left[\mathbb{E}[Y^{1}\mid A=1,\mathbf{X}_{\tau}\cup\mathbf{X}_{S},S=\mathrm{e}]\mid\mathbf{X}_{\tau}\right]-\mathbb{E}\!\left[\mathbb{E}[Y^{0}\mid A=0,\mathbf{X}_{\tau}\cup\mathbf{X}_{S},S=\mathrm{e}]\mid\mathbf{X}_{\tau}\right]$
(linearity of conditional expectation)
$=\mathbb{E}\!\left[\mathbb{E}[Y\mid A=1,\mathbf{X}_{\tau}\cup\mathbf{X}_{S},S=\mathrm{e}]\mid\mathbf{X}_{\tau}\right]-\mathbb{E}\!\left[\mathbb{E}[Y\mid A=0,\mathbf{X}_{\tau}\cup\mathbf{X}_{S},S=\mathrm{e}]\mid\mathbf{X}_{\tau}\right]$
(Assumption A1)
$=\mathbb{E}[Y\mid A=1,\mathbf{X}_{\tau},S=\mathrm{e}]-\mathbb{E}[Y\mid A=0,\mathbf{X}_{\tau},S=\mathrm{e}]$
(tower property)
Appendix B Illustrative example
@@
Figure 3: Comparative illustration between Independent GPs and Multitask GP. Here, AA represents distinct treatment groups, $Y$ corresponds to the outcome, and $X$ denotes baseline covariates.
Appendix C Derivation of Posterior Distribution
@@
$\mathbb{E}\left[y^{\text{e}}_{i}y^{\text{e}}_{j}\right]$ $=\begin{cases}\beta^{\text{e}}k(\mathbf{x}^{\text{e}}_{i},\mathbf{x}^{\text{e}}_{j})+\sigma^{2}\quad&\text{if }i=j\\ \beta^{\text{e}}k(\mathbf{x}^{\text{e}}_{i},\mathbf{x}^{\text{e}}_{j})\quad&\text{otherwise}\end{cases}$
@@
$\mathbb{E}\left[y^{\text{o}}_{i}y^{\text{o}}_{j}\right]$ $=\begin{cases}\beta^{\text{o}}k(\mathbf{x}^{\text{o}}_{i},\mathbf{x}^{\text{o}}_{j})+\sigma^{2}\quad&\text{if }i=j\\ \beta^{\text{o}}k(\mathbf{x}^{\text{o}}_{i},\mathbf{x}^{\text{o}}_{j})\quad&\text{otherwise}\end{cases}$
@@
[…]
@@
Appendix D Interpreting $\rho$
Proposition D.1.
We have $\rho=1$ if and only if
Proof D.2.
It is easier to work with $1-\rho^{2}$, where plugging in the values from (Equation 1) gives
The denominator is positive and finite, so the numerator is thus 0 if and only if $\alpha_{1}^{\text{e}}\alpha_{2}^{\text{o}}=\alpha_{1}^{\text{o}}\alpha_{2}^{\text{e}}$.
Appendix E Proof of Proposition 3.1
@@ Appendix D Interpreting $\rho$
$B^{\prime}=\begin{bmatrix}(\beta^{\text{e}})^{2}&\beta^{\text{e}}\beta^{\text{eo}}\\ \beta^{\text{e}}\beta^{\text{eo}}&(\beta^{\text{eo}})^{2}\end{bmatrix},\penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ \penalty 10000\ k^{\prime}(\mathbf{x}_{*},\mathbf{x}_{*})=k(\mathbf{x}_{*},X^{\text{e}})^{T}{\Sigma_{\text{e}}}^{-1}k(\mathbf{x}_{*},X^{\text{e}})$
@@
Appendix F Confounding function